OpenStudy (anonymous):

Consider the following optimization problem: optimize z(x, y) = 2xy subject to x^2 + xy + 4y^2 = 40 x ≥ 0 and y ≥ 0. This is an exam question and i did it wrong. basically, the Lagrange criteria says this is not feasible but i dont understand why

OpenStudy (freckles):

\[z=2xy \\ x^2+xy+4y^2=40 \\z_x=\lambda g_x \\ z_y=\lambda g_y \\ x^2+xy+4y^2=40\] so we want to solve the bottom 3 equations like I'm just trying to see if it does look not feasible

OpenStudy (freckles):

\[2y= \lambda (2x+y) \\ 2x=\lambda(x+8y) \\ x^2+xy+4y^2=40\] so you played with these 3 equations right?

OpenStudy (freckles):

I actually think it works if I didn't make a mistake.

OpenStudy (freckles):

I'm sure you got somewhere that \[8y^2=2x^2 \\ \text{ or } 4y^2=x^2 \] plug that into your constraint

OpenStudy (freckles):

are you there?

OpenStudy (freckles):

this also might help since y>=0 then y|y| can be written as y^2

OpenStudy (anonymous):

yes i got that, but could u explain why the points are not feasible to start with?

OpenStudy (freckles):

What do you mean by feasible? I was able to obtain point through lagrange

OpenStudy (freckles):

\[4y^2+2y^2+4y^2=40 \text{solve for y then x comes pretty easily afterwards} \]